OpenStudy (anonymous):

Solve the following system of equations: 2x - 5y =22 -2x - 3y = 6 Answer a. (1, -8) b. (-9, -8) c. (-9, 3) d. (8, 9)

OpenStudy (anonymous):

THERE IS A PROBLEM WITH THE QUESTION..

OpenStudy (anonymous):

Systems Of Equations

OpenStudy (phi):

you must have a typo in the equations.

OpenStudy (phi):

are you sure the second equation has -2x ?? how about 2x - 3y = 6 instead?

OpenStudy (anonymous):

no on the thing its a -

OpenStudy (phi):

In that case you are suffering from "bogus question" syndrome. because the solution to the system you posted is x= 9/4 and y= -7/2

OpenStudy (anonymous):

maby take out the - then when its solve plug it in?

OpenStudy (phi):

on the other hand, the solution to 2x - 5y =22 2x - 3y = 6 is (-9,-8) so maybe there is a typo in the question.

OpenStudy (phi):

maby take out the - then when its solve plug it in? no, that isn't allowed.

OpenStudy (anonymous):

yeah i was thinking that i enter you anser see if it was a typo.

OpenStudy (phi):

If you have a teacher, I would point this one out to them

OpenStudy (phi):

and it would be good if you could solve the original problem. tho none of the answers match, it is good practice, and it proves you can solve these things.